OpenStudy (lgbasallote):

\[\int \frac{(4v^2 + 4v + 1)dv}{v^3 + v^2}\] how?? o.O partial fractions??

OpenStudy (shubhamsrg):

assuming denom to be v^2 + v its integral (4 + 1/(v^2 + v) dv) shouldnt be much tough..denom is (v)(v+1)

sam (.sam.):

\[\int\limits \frac{1+4 v+4 v^2}{v^3+v^2} \, dv\] Split \[\int\limits \frac{4 v^2+4 v+1}{v^2 (v+1)} \, dv\] Partial fraction

OpenStudy (anonymous):

More specifically, \[\frac{4v^2+4v+1}{v^3+v^2}=\frac{4v^2+4v+1}{v^2(v+1)}=\frac{Av+B}{v^2}+\frac{C}{v+1}\]

OpenStudy (lgbasallote):

i figured it out lol sorry

OpenStudy (shubhamsrg):

ohh am sorry i thought you by mistake wrote v^2+ v^2 never noticed it to be v^3.. hmmn so bottom line is : you know any good eye specialist? :P